Database last updated on Thu, 29 Jan 2026 16:10:51 +0000 by localhost
Instrument constant changes for the week before 2024300 (2 sec, 0 rows)
SL burned out? (AP results when SL is ON) (0 sec, 3 rows)
HG burned out? (AP results when HG is ON) (0 sec, 1 rows)
Measurements and tests summary.
| _jday | _loc | _serial |
| 300 | TORONTO | #254 |
(0 sec)
Number of days since last dome/window cleaning (0 sec, 1 rows)
Run/Stop results for last week. (0 sec, 1 rows)
| Serial | Total | Failed 0 | Failed dark | Failed 1 | Failed 2 | Failed 3 | Failed 4 | Failed 5 | Failed double |
| 254 | 13 | 0 | 1 | 0 | 0 | 0 | 0 | 0 | 0 |
Ozone results for good DS obs (std<2.5DU, mu<4) within 3 minutes between 2 Brewers. (0 sec)
No ozone measurements to display.
Unsorted comments (1 sec, 11872 rows)
The END (0 sec, 2 rows)